# Intelligence Test

### v1

@@ 7 Conclusions and Future Work
In this paper, we introduce Intelligence Test. It is inspired by Natural Selection and quantifies the intelligence of any subject in any task. It demonstrates three advantages.
@@
  • Secondly, Intelligence Test provides a roadmap for the future development of AI. It shows a clear relationship to the scale of AI systems, which enables us to project the time to reach high-level intelligence.
  • Finally, Intelligence Test helps reveal the nature of human tasks and AI. It suggests that human tasks exhibit criticality properties and AI’s superficial manner to solve human tasks.
We believe that future efforts should focus on three key areas. Firstly, since current AI technologies are mostly at the Limited Level, we need to identify appropriate application scenarios and design an effective human supervision framework. Secondly, since Intelligence Test connects the model size to the intelligence levels, we can use Intelligence Test to plan the roadmap of future AI development. Finally, linking Intelligence Test to SOC theory shows promising results and more efforts shall be made in this direction. This will not only help design more effective AI models but also deepen our understanding of humans ourselves.
@@
Jingtao Zhan initiated the project, proposed Intelligence Test, conducted experiments, and used SOC for theoretical analysis. Jiahao Zhao contributed to the experiments by conducting the text search task. Jiayu Li contributed to the experiments by conducting the recommendation system task. Yiqun Liu supervised the project, engaged in multiple discussions, and guided the researchers in exploring the fundamental principles underlying the observed phenomena. Bo Zhang participated in multiple discussions, provided important advice, and pointed out the similarity between the experimental results and phenomenon in SOC systems. Qingyao Ai, Jiaxin Mao, Hongning Wang, Min Zhang, and Shaoping Ma were involved in the early-stage discussions and provided invaluable feedback. Jingtao Zhan drafted the manuscript, and all other authors contributed important insights and suggestions to improve the writing.

### v2

@@ 7 Conclusions and Future Work
In this paper, we introduce Survival Game. It is inspired by Natural Selection and quantifies the intelligence of any subject in any task. It demonstrates three advantages.
@@
  • Secondly, Survival Game provides a roadmap for the future development of AI. It shows a clear relationship to the scale of AI systems, which enables us to project the time to reach high-level intelligence.
  • Finally, Survival Game helps reveal the nature of human tasks and AI. It suggests that human tasks exhibit criticality properties and AI’s superficial manner to solve human tasks.
We believe that future efforts should focus on three key areas. Firstly, since current AI technologies are mostly at the Limited Level, we need to identify appropriate application scenarios and design an effective human supervision framework. Secondly, since Survival Game connects the model size to the intelligence levels, we can use Survival Game to plan the roadmap of future AI development. Finally, linking Survival Game to SOC theory shows promising results and more efforts shall be made in this direction. This will not only help design more effective AI models but also deepen our understanding of humans ourselves.
@@
Jingtao Zhan initiated the project, proposed Survival Game, conducted experiments, and used SOC for theoretical analysis. Jiahao Zhao contributed to the experiments by conducting the text search task. Jiayu Li contributed to the experiments by conducting the recommendation system task. Yiqun Liu supervised the project, engaged in multiple discussions, and guided the researchers in exploring the fundamental principles underlying the observed phenomena. Bo Zhang participated in multiple discussions, provided important advice, and pointed out the similarity between the experimental results and phenomenon in SOC systems. Qingyao Ai, Jiaxin Mao, Hongning Wang, Min Zhang, and Shaoping Ma were involved in the early-stage discussions and provided invaluable feedback. Jingtao Zhan drafted the manuscript, and all other authors contributed important insights and suggestions to improve the writing.